There wewe 723 babies born last december there were 215 more girls than boys. How many of each were born?
Answer by checkley77(12844)

x+x+215=723
2x=723-215
2x=508
x=508/2
x=254 boys.
254+215=469 girls.
Proof
254+469=723
723=723